2.5 g of ammonium chloride and 5.02 g of powdered hexaaquacobalt(II) chloride (CoCl2 * 6 H2O) dissolved in 15 mL of 15M aqueous ammonia.

4 mL of 30% hydrogen peroxide is added from a dropper pipette. 2.5 g of sodium nitrite added to the solution while stirring to dissolve.

What is theoretical yield of cobalt complex?
Answer #1

The complex formed in this synthesis is (Co(NH3)6)Cl3

Molar mass of the complex is 267.4 g/mole

In this synthesis limiting reagent is Cobalt chloride.

Molar mass of cobalt chloride 237.923 g/mole

Mass of cobalt chloride = 5.02/237.923 = 0.021 moles.

Thereotical yield = moles of cobalt chloride*molar mass of product

= 0.021*267.4 = 5.6154 grams.